OpenStudy (anonymous):

Will FAN!!! What is -12 divided by 2/3rds?

OpenStudy (anonymous):

Do you know how to set it up?

OpenStudy (qwertty123):

@Mc2 You there?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

sorry really slow computer

OpenStudy (anonymous):

yes im here

OpenStudy (anonymous):

ok i think i set it up like this -12/1 divided by 3/2=

OpenStudy (anonymous):

Hello?

OpenStudy (qwertty123):

okay

OpenStudy (qwertty123):

yes hold on my computer is so slow!

OpenStudy (qwertty123):

are we simplifying this?

OpenStudy (qwertty123):

@Mc2

OpenStudy (mrnood):

when you divide by a fraction it is the same as multiplying by the 'inverse' of the fraction - that is 'turn the fraction oer and then multiply so 12 divided by 3/4 is the same as 12 TIMES 4/3 Do the same with your question and the answer will be clear..

OpenStudy (mathmale):

Good, EXCEPT that we should be "multiplying by the 'RECIPROCAL' of the fraction - that is 'turn the fraction oer and then multiply" "inverse" has a different meaning in algebra. y=ln x and y=e^x are "inverse functions."

OpenStudy (mathmale):

correctly expressed symbolically, -21 divided by two thirds is\[\frac{ -21 }{ \frac{ 2 }{ 3 } }\]

OpenStudy (mathmale):

Invert the fraction in the denominator and then multiply the result by -21.